# Cloud2BIM: An open-source automatic pipeline for efficient conversion of large-scale point clouds into IFC format

### v1

@@ Abstract
Building Information Modeling (BIM) is an essential component in the sustainable reconstruction and revitalization of ageing structures. However, model creation usually relies on laborious manual transformation of the unstructured point cloud data provided by laser scans or photogrammetry. This paper presents Cloud2BIM, an open-source software tool designed to automate the conversion of point clouds into BIM models compliant with the Industry Foundation Classes (IFC) standard. Cloud2BIM integrates advanced algorithms for wall and slab segmentation, opening detection, and room zoning based on real wall surfaces, resulting in a comprehensive and fully automated workflow. Unlike existing tools, it avoids computationally- and calibration-intensive techniques such as RANSAC, supports non-orthogonal geometries, and provides unprecedented processing speed—achieving results up to seven times faster than fastest competing solutions. Systematic validation using benchmark datasets confirms that Cloud2BIM is an easy-to-use, efficient, and scalable solution for generating accurate BIM models, capable of converting extensive point cloud datasets for entire buildings into IFC format with minimal user input.
@@ 1 Introduction
Over the last decade, the construction industry has experienced a profound digital transformation, marking a new architectural and engineering design epoch. This period, commonly referred to as Construction 4.0 (Sawhney et al., 2020), is characterized by a pivotal shift toward the utilization of advanced digital technologies aimed at enhancing the efficiency, quality, safety, and sustainability of construction projects (Statsenko et al., 2023).
Amid these advancements, there is a growing focus on sustainability, especially for buildings nearing the end of their lifecycle. This focus has spurred construction companies to develop and adopt innovative methods that significantly reduce the environmental impacts of building and demolition activities. In Europe, construction and demolition account for approximately 36% of total waste generation, highlighting a considerable environmental challenge (Eurostat, 2020).
In response to the demand for sustainable construction practices, the industry is increasingly retrofitting and modernizing existing buildings. These initiatives typically involve upgrading insulation, installing energy-efficient fixtures, and integrating renewable energy systems to enhance building performance and sustainability (Dakwale et al., 2011; Grecchi, 2022). Furthermore, adopting deconstruction and recycling practices provides a practical alternative by reducing waste and decreasing the dependence on new, raw materials (Ge et al., 2017). Although deconstruction can be slower and more costly than traditional demolition (Bertino et al., 2021), it is crucial for transitioning the construction industry towards a circular economy.
However, the effective implementation of these sustainable strategies is often hampered by the absence or obsolescence of building documentation. Building models, essential for documentation, can be accurately created using laser scanning or photogrammetry. Both methods generate detailed three-dimensional coordinates and potentially color information of points on the surfaces of structures. The density of these point clouds varies depending on the technology used and the distance from which the scans are taken. Photogrammetry employs conventional RGB images to construct point clouds (Aber et al., 2019), whereas laser scanning utilizes laser beams to capture comprehensive 3D data, forming detailed point clouds (Valero et al., 2022).
Moreover, as Mirzaei et al. (2023) report, these point clouds are invaluable for extracting precise geometry information about individual building members, aiding in their dimensional quality control during construction or renovation projects. Despite their high accuracy and visually precise representations, point clouds are inherently unstructured and do not support database operations such as listing individual elements, nor do they contain contextual information. In contrast, entity-based representations like building information models (BIM) incorporate detailed metadata and relational data within each component, enabling thorough analysis and manipulation. The capability to transform point clouds into precise, digital, entity-based models facilitates the planning of retrofits or reconstructions with enhanced accuracy, supporting conservation and sustainable development goals (Akbarieh et al., 2020).
Converting point cloud data into usable BIM, or scan-to-BIM process, is often a meticulous and labor-intensive task that demands significant expertise and advanced software (Klinc et al., 2021; Abreu et al., 2023). This procedure is predominantly managed by skilled BIM modelers manually integrating point cloud data with parametric objects. While this manual approach is practical, it is also highly time-consuming and results in considerable costs. This manual labor requirement presents a challenge, particularly for small investors, as the high costs of manually created 3D models and the subsequent generation of BIM are often prohibitive despite the potential value of such models. Recent developments have introduced both semi-automatic (Dore and Murphy, 2014) and fully automatic (Kwadjo et al., 2019; Romero-Jarén and Arranz, 2021; Wang et al., 2015; Martens and Blankenbach, 2023; Mahmoud et al., 2024) solutions to address this issue.
The available scan-to-BIM solutions exploit various algorithms for segmenting and classifying building elements. For segmenting slabs, analyzing point density along the vertical axis using height histograms, as explored in a few studies (Oesau et al., 2014; Turner and Zakhor, 2014; Romero-Jarén and Arranz, 2021; Martens and Blankenbach, 2023), became a standard method. This approach is simple yet sensitive to noise, which can lead to inaccuracies. As an alternative, Nikoohemat et al. (2018) introduced an algorithm that detects horizontal planes by analyzing trajectories from mobile laser scanners, offering effective results but requiring specific scanning technology.
Wall classification often employs the random sample consensus (RANSAC) algorithm, which identifies sets of points that align with a plane, frequently utilized for detecting interior walls (Schnabel et al., 2007; Ochmann et al., 2016; Murali et al., 2017; Ochmann et al., 2019). Another common strategy simplifies scenes to a 2D plan view, analyzing wall positions under the assumption that walls are vertical, which simplifies processing but restricts the geometrical possibilities (Oesau et al., 2014; Bormann et al., 2016; Macher et al., 2017; Jung et al., 2018).
Results from wall geometry analysis can be categorized into two main approaches: (i) room-based, which focuses on the dimensions of spaces and rooms (Wang et al., 2017; Yang et al., 2019), and (ii) wall-based or volumetric, which concentrates on wall geometry and the 3D model consisting of wall objects (Ochmann et al., 2019; Nikoohemat et al., 2018; Bassier and Vergauwen, 2020; Martens and Blankenbach, 2023; Jung et al., 2018). The room-based approach typically identifies only the surfaces of rooms, whereas the wall-based approach involves identifying opposing surfaces that form walls and solving their topology and connections.
Despite the plethora of semi-automated and automated scan-to-BIM solutions available, a fully open-source, universally applicable tool independent of commercial software is still lacking in the current literature. While some initiatives have made source code available, such as the notable work of Bassier and Vergauwen (2020), these examples are the exception rather than the rule. The lack of widely available code hinders the ability of the construction industry to adopt and adapt new technologies and methodologies effectively. In addition, many existing solutions rely on proprietary software to function, such as the Dynamo add-on [^1] https://dynamobim.org/ with Revit software by Chen (2022) and Mahmoud et al. (2024), or depend on the Revit API as documented by Mehranfar et al. (2024b).
This paper introduces Cloud2BIM, a novel Python-based open-source tool that fills the gap in the current scan-to-BIM landscape. It provides a fully automated and flexible conversion process, free from dependencies on commercial software. Unlike existing methods, Cloud2BIM handles complex architectural geometries, such as non-orthogonal wall configurations, and efficiently processes large data sets with minimal user input. A distinctive feature of Cloud2BIM is its ability to export data directly in IFC format, ensuring interoperability across different BIM software platforms. It also includes advanced zoning capabilities that more accurately define space characteristics using actual wall surfaces. By simplifying the conversion process and offering it as an open-source solution, Cloud2BIM makes advanced BIM technology available to a broader audience, promoting sustainable construction practices through precise and rapid transformation of raw scan data into usable BIM outputs.
2 Testing datasets
The process of data dilution was executed using the open-source CloudCompare software, where the point clouds were diluted to a prescribed minimum distance to the nearest point ($d_{\mathrm{min}}$). In ’spatial’ mode, the user specifies a minimum distance between points. CloudCompare then selects points from the original cloud, ensuring no point in the resulting cloud is closer to another than the specified distance. As this value increases, fewer points are retained. The underlying algorithm operates as follows: Points are processed in an order determined by their spatial organization within an octree structure. The initial point is marked as ’to be kept’. For each subsequent point, the algorithm checks for neighbors within the specified radius. If no neighboring points are marked as ’to be kept’, this point is also marked. Ultimately, only points marked as ’to be kept’ are included in the final output cloud. The impact of this dilution is illustrated in Figure 1.
Such an approach to dilution of point clouds proved effective for buildings with relatively simple geometric details. However, for buildings of historical significance that feature intricate decorations, the dilution strategy proposed by Štroner et al. (2022) is more suitable. This method adjusts the dilution factor based on the local curvature, helping to preserve critical features while reducing the risk of feature loss, often referred to as ’feature washing.’ Caution is advised with this approach as it tends to remove points from planar surfaces, which could lead to insufficient density for effective point cloud segmentation. Although avoiding dilution prevents data loss, it dramatically increases computational demands and may not always be practical.
2.1 ETH Zürich synth3
The synthetic dataset, provided by the Multimedia Laboratory at the University of Zurich222www.ifi.uzh.ch/en/vmml/research/datasets.html, offers a complex challenge for testing the construction of volumetric walls under the Atlanta-world assumption (Table 1). This assumption permits the depiction of scenes using vertical and horizontal planes in 3D, without the constraint of vertical planes (walls) being orthogonal to each other (Pintore et al., 2020). The dataset encompasses 20 interior walls, features 13 door openings, and is designed for algorithms that reconstruct interior walls connecting at various angles. Notably, the external surfaces of the object were not created by the dataset authors, simulating purely indoor scanning. This dataset has been pivotal in research focused on indoor wall detection, contributing significantly to other studies (Ochmann et al., 2019; Shi et al., 2019; Bassier and Vergauwen, 2020).
Dataset name Visualization $n_{\mathrm{points}}$ Size $d_{\mathrm{min}}$
ETH Zürich synth3 [Uncaptioned image] 6,905,684 468 kB 0.01
Table 1: ETH Zürich synth3 dataset specification.
2.2 Kladno station
Point cloud data from Kladno railway station (Table 2), captured using the terrestrial laser scanner, was employed to demonstrate the wall segmentation capabilities of the developed software on a real-world dataset. For this purpose, one floor of the building was utilized. Using an actual point cloud from this building provides the benefit of including real-world irregularities and imperfections that the program must navigate, thereby enhancing the robustness and applicability of the segmentation results.
Dataset name Visualization $n_{\mathrm{points}}$ Size $d_{\mathrm{min}}$
Kladno station [Uncaptioned image] 2,551,721 190 kB 0.025
Table 2: Kladno station dataset specification.
2.3 Prague hotel
Dataset name Visualization $n_{\mathrm{points}}$ Size $d_{\mathrm{min}}$
Prague hotel [Uncaptioned image] 40,152,849 2.54 GB 0.025
Table 3: Prague hotel dataset specification.
@@ 3 Methods
This section outlines the algorithms used to obtain the geometric representation of building structures through Cloud2BIM, our Python-based software that adheres to open-source principles. A comprehensive list of the program dependencies and packages utilized during development is provided in Table 4. The entire codebase is accessible on our GitHub repository. Figure 2 presents a diagram illustrating the program workflow.
Figure 2: Flowchart illustrating the process of generating a 3D model of a building from laser scans, detailing each step from data acquisition to model completion.
@@ 3.1 Identification of slabs
Cloud2BIM adopts a widely used method that analyzes point density within a histogram to identify horizontal planes in point clouds (Bassier and Vergauwen, 2020), illustrated in Figure 3. This approach assumes that the floor slab is horizontal relative to the $x$-$y$ plane, has uniform thickness, and lacks local height variations like steps. A candidate for a horizontal surface is identified within a height range that contains at least 50% of the maximum number of points found in any single plane. This threshold is adjustable to account for uneven point distribution often seen in real-world scans, where fewer points may be recorded at floor level due to obstructions like furniture or scanner positioning.
@@
Using the algorithm by Suzuki and Abe (1985) implemented through OpenCV’s findContours() function, the most prominent contour is extracted, scaled, and converted into a Polygon object that delineates the floor plan. To simplify this contour, the Douglas-Peucker algorithm (Douglas and Peucker, 1973) is applied, reducing the number of points and smoothing the shape, as shown in Figure 6. This process effectively represents the polygon of a slab with fewer, more significant points. The complete program workflow for slab segmentation is depicted in Figure 7.
@@
Figure 7: Flowchart diagram illustrating the function used to detect slabs in an unorganized point cloud.
@@ 3.2 Segmentation into storeys
Figure 8: Flowchart diagram illustrating the process of segmenting a point cloud into individual storeys, highlighting the sequence of operations from slab identification to storey-specific point cloud extraction.
@@ 3.3 Identification of walls
Two primary strategies prevail for reconstructing interior walls from point cloud data: room-based and wall-based. The room-based approach simplifies the segmentation process by focusing on reconstructing the geometry of individual rooms, which involves segmenting the point cloud into distinct spaces and determining the optimal shape for each. This approach, however, struggles with flexible parametric wall geometry and unenclosed spaces. On the other hand, the wall-based approach clusters point cloud data around each wall, aiming to reconstruct topologically consistent structures (Bassier and Vergauwen, 2020).
The Cloud2BIM’s Identify_walls function, described using the flowchart in Figure 9, detects and defines wall geometry within a point cloud. It isolates individual wall segments within each floor’s horizontal section, set by default to 90–100%
Figure 9: Flowchart diagram illustrating the algorithms implemented in the Identify_walls function for segmentation of walls.
This 2D point cloud then undergoes histogram analysis to create a 2D histogram, where each bin’s value reflects the point density in that spatial region. Converting this histogram into a binary image based on a predefined threshold isolates regions with high point densities (likely indicative of walls). This binary image (Figure 10) serves as the basis for further wall detection and geometry definition. It undergoes processing to extract contours using the findContours function from the OpenCV library (Bradski, 2000), explicitly targeting the outermost boundaries of wall segments. Each detected contour is then approximated into line segments through the Douglas-Peucker algorithm. These segments are converted from pixel-based coordinates to real-world coordinates.

### v2

@@ Abstract
Building Information Model (BIM) creation usually relies on laborious manual transformation of the unstructured point cloud data provided by laser scans or photogrammetry. This paper presents Cloud2BIM, an open-source software tool designed to automate the conversion of point clouds into BIM models compliant with the Industry Foundation Classes (IFC) standard. Cloud2BIM integrates advanced algorithms for wall and slab segmentation, opening detection, and room zoning based on real wall surfaces, resulting in a comprehensive and fully automated workflow. Unlike existing tools, it avoids computationally- and calibration-intensive techniques such as RANSAC, supports non-orthogonal geometries, and provides unprecedented processing speed, achieving results up to seven times faster than fastest competing solutions. Systematic validation using benchmark datasets confirms that Cloud2BIM is an easy-to-use, efficient, and scalable solution for generating accurate BIM models, capable of converting extensive point cloud datasets for entire buildings into IFC format with minimal user input.
@@ 1 Introduction
The construction industry has undergone a rapid digital transformation over the past decade, ushering in a new era of architectural and engineering design. This transformation, commonly called Construction 4.0 [1], is driven by advancements in digital technologies that improve efficiency, quality, safety, and interoperability in construction projects [2]. A key aspect of this transformation is the increasing reliance on digital workflows to document and manage the lifecycle of buildings. However, these workflows remain challenging to implement in existing buildings due to the lack of reliable digital documentation.
The absence or obsolescence of building documentation presents a significant barrier to the efficient retrofitting and maintenance of existing structures. Accurately representing buildings in digital form is crucial for various applications, including energy efficiency upgrades, structural assessments, and facility management. Point cloud-based workflows are increasingly employed to capture detailed three-dimensional representations of buildings to address this. Laser scanning and photogrammetry are the primary technologies used for this purpose, generating high-density point clouds that record the spatial geometry of building elements with high precision [3, 4].
Despite the high accuracy of point cloud data, raw point clouds lack semantic structure and are unsuitable for direct use in most engineering and architectural applications. Building information modeling (BIM) provides a structured digital representation of a building, integrating geometric and semantic information into a single framework. Converting point clouds into BIM, commonly referred to as scan-to-BIM, usually remains a labor-intensive task that relies heavily on manual input from skilled professionals [5].
Recent research efforts have sought to reduce the manual workload associated with scan-to-BIM by introducing methods providing various levels of automation. Semi-automatic approaches typically assist users in identifying structural elements but still require significant manual intervention [6]. Fully automated solutions aim to extract building components without user input, leveraging rule-based segmentation, deep learning, and geometric fitting techniques [7, 8, 9, 10, 11]. While these methods have demonstrated promising results, they often suffer from computational inefficiencies, limitations in handling complex geometries, or dependencies on specific commercial software environments.
The (semi-)automatic scan-to-BIM tools employ a combination of computational techniques to segment and classify building elements within point clouds. These include point cloud density histograms along selected axes [12, 13, 8, 10], trajectory-based plane detection that relies on specialized mobile laser scanning systems [14], and deep learning models designed for structural element segmentation [15], though these primarily focus on classification rather than generating complete BIM-compatible geometries. Wall detection methods typically follow either a room-based or wall-based approach. Room-based methods identify enclosed spaces and approximate their boundaries [16, 17], but often struggle with incomplete or missing data. In contrast, wall-based techniques reconstruct individual walls by detecting their surfaces and resolving spatial relationships [18, 14, 19, 10, 20], offering more precise 3D representations at the cost of additional preprocessing. A widely used strategy for extracting walls from point clouds involves the RANSAC algorithm, which detects planar surfaces [21, 22, 23, 18, 24, 25], but requires careful parameter tuning and is sensitive to cluttered indoor environments. An alternative approach simplifies the scene into a 2D plan view, assuming vertical walls aligned to a regular grid [12, 26, 27, 20], improving processing speed but limiting adaptability to irregular or non-standard building geometries.
Based on the authors’ experience, existing scan-to-BIM approaches face critical drawbacks that hinder broader adoption. Many solutions struggle with parameter sensitivity, complex indoor environments, or the need for extensive pre-processing, making them inefficient for handling diverse architectural styles. Additionally, a significant portion of available tools depends on proprietary software ecosystems [28, 11, 29, 30], restricting adaptability and limiting accessibility for a broader audience. While some projects have made their code publicly available [19], most remain tied to commercial platforms, reducing transparency and customization options. This reliance on closed ecosystems prevents researchers and practitioners from refining or adapting workflows to specific project needs, ultimately slowing innovation in the field. Furthermore, some seemingly promising approaches operate with significant constraints, such as reconstructing only visible surfaces rather than full volumetric elements [31].
This paper introduces Cloud2BIM, a fully open-source, Python-based scan-to-BIM tool that automates the conversion of large-scale point clouds into industry foundation classes (IFC)-compliant BIM models without relying on proprietary software. It features a fully automated processing pipeline that efficiently segments and classifies structural elements, including slabs, walls, and openings, while accurately handling non-orthogonal geometries. Unlike many existing methods, Cloud2BIM reconstructs building topology explicitly, ensuring that individual components are correctly structured. The direct IFC export enables seamless interoperability with BIM platforms such as Revit, Archicad, and FreeCAD, while also allowing compatibility with various free and open-source software tools for IFC editing, including Blender with the Bonsai plugin. Additionally, it integrates zoning capabilities, deriving room boundaries from actual wall surfaces for improved spatial accuracy in applications such as energy modeling or facility management. By offering a computationally efficient, scalable, and fully transparent workflow, Cloud2BIM removes barriers to adoption, allowing users to refine and adapt scan-to-BIM processes to diverse architectural and engineering needs.
@@ 2 Methods
This section outlines the algorithms used to obtain the geometric representation of building structures (Figure 1) through Cloud2BIM, our Python-based software that adheres to open-source principles. A comprehensive list of the program dependencies and packages utilized during development is provided in Table 1. The entire codebase is accessible on our GitHub repository [32]. Figure 2 presents a diagram illustrating the program workflow.
Figure 1: Overview of the Cloud2BIM workflow, illustrating the automated conversion from raw point clouds to IFC BIM models, including detection of slabs, walls, zones, and openings, and their conversion to the IFC format.
Figure 2: Flowchart illustrating the process of generating a 3D model of a building from laser scans, detailing each step from data acquisition to model completion. The implementation of this workflow can be found in the cloud2entities.py file [32].
@@
2.1 Testing datasets
The process of data dilution was executed using the open-source CloudCompare software, where the point clouds were diluted to a prescribed minimum distance to the nearest point ($d_{\mathrm{min}}$). In ’spatial’ mode, the user specifies a minimum distance between points. CloudCompare then selects points from the original cloud, ensuring no point in the resulting cloud is closer to another than the specified distance. As this value increases, fewer points are retained. The underlying algorithm operates as follows: Points are processed in an order determined by their spatial organization within an octree structure. The initial point is marked as ‘to be kept’. For each subsequent point, the algorithm checks for neighbors within the specified radius. If no neighboring points are marked as ’to be kept’, this point is also marked. Ultimately, only points marked as ’to be kept’ are included in the final output cloud. The impact of this dilution is illustrated in Figure 3.
Such an approach to dilution of point clouds proved effective for buildings with relatively simple geometric details. However, for buildings of historical significance that feature intricate decorations, the dilution strategy proposed by [33] is more suitable. This method adjusts the dilution factor based on the local curvature, helping to preserve critical features while reducing the risk of feature loss, often referred to as ’feature washing.’ Caution is advised with this approach as it tends to remove points from planar surfaces, which could lead to insufficient density for effective point cloud segmentation. Although avoiding dilution prevents data loss, it dramatically increases computational demands and may not always be practical.
2.2 ETH Zürich synth3
The synthetic dataset, provided by the Multimedia Laboratory at the University of Zurich111www.ifi.uzh.ch/en/vmml/research/datasets.html, serves as a challenging benchmark for volumetric wall reconstruction. It is designed to evaluate methods that operate under the Atlanta-world assumption, which allows for non-orthogonal wall configurations (Table 2). This assumption permits the depiction of scenes using vertical and horizontal planes in 3D, without the constraint of vertical planes (walls) being orthogonal to each other [34] as assumed, e.g., by Kong et al. [35] or Kim and Lee [36] in their recent studies on classification of point clouds. The dataset encompasses 20 interior walls, features 13 door openings, and is designed for algorithms that reconstruct interior walls connecting at various angles. Notably, the external surfaces of the object were not created by the dataset authors, simulating purely indoor scanning. This dataset has been pivotal in research focused on indoor wall detection, contributing significantly to other studies [18, 37, 19].
Dataset name Visualization $n_{\mathrm{points}}$ Size $d_{\mathrm{min}}$
ETH Zürich synth3 [Uncaptioned image] 6,905,684 468 kB 0.01
Table 2: ETH Zürich synth3 dataset specification.
2.3 Kladno station
Point cloud data from Kladno railway station (Table 3), captured using the terrestrial laser scanner, was employed to demonstrate the wall segmentation capabilities of the developed software on a real-world dataset. For this purpose, one floor of the building was utilized. Using an actual point cloud from this building provides the benefit of including real-world irregularities and imperfections that the program must navigate, thereby enhancing the robustness and applicability of the segmentation results. The complete original point cloud is available at Zenodo platform  [^2] https://zenodo.org/records/14221915.
Dataset name Visualization $n_{\mathrm{points}}$ Size $d_{\mathrm{min}}$
Kladno station [Uncaptioned image] 2,551,721 190 kB 0.025
Table 3: Kladno station dataset specification.
2.4 Prague hotel
Dataset name Visualization $n_{\mathrm{points}}$ Size $d_{\mathrm{min}}$
Prague hotel [Uncaptioned image] 40,152,849 2.54 GB 0.025
Table 4: Prague hotel dataset specification.
@@ 2.5 Identification of slabs
Cloud2BIM adopts a widely used method that analyzes point density within a histogram to identify horizontal planes in point clouds [19], illustrated in Figure 4. This approach assumes that the floor slab is horizontal relative to the $x$-$y$ plane, has uniform thickness, and lacks local height variations like steps. A candidate for a horizontal surface is identified within a height range that contains at least 50% of the maximum number of points found in any single plane. This threshold is adjustable to account for uneven point distribution often seen in real-world scans, where fewer points may be recorded at floor level due to obstructions like furniture or scanner positioning. Similar approach was used e.g. by Chen et al. [38], who estimated stiffness of structures based on point cloud density distribution in order to assess their performance during earthquakes.
@@
Using the algorithm by Suzuki and Abe [39] implemented through OpenCV’s findContours() function, the most prominent contour is extracted, scaled, and converted into a Polygon object that delineates the floor plan. To simplify this contour, the Douglas-Peucker algorithm [40, 41] is applied, reducing the number of points and smoothing the shape, as shown in Figure 7. The algorithm iteratively reduces the number of points defining a polyline and operates by recursively eliminating points that deviate less than a user-defined tolerance, epsilon, from the simplified line segments. This process effectively represents the polygon of a slab with fewer, more significant points. The complete program workflow for slab segmentation is depicted in Figure 8.
@@
Figure 8: Flowchart diagram illustrating the function used to detect slabs in an unorganized point cloud. The implementation of this workflow can be found in the aux_functions.py file, function indentify_slabs, lines 196–307 [32].
@@ 2.6 Segmentation into storeys
Figure 9: Flowchart diagram illustrating the process of segmenting a point cloud into individual storeys, highlighting the sequence of operations from slab identification to storey-specific point cloud extraction. The implementation of this workflow can be found in the aux_functions.py file, function split_pointcloud_to_storeys, lines 310–328 [32].
@@ 2.7 Identification of walls
Two primary strategies prevail for reconstructing interior walls from point cloud data: room-based and wall-based. The room-based approach simplifies the segmentation process by focusing on reconstructing the geometry of individual rooms, which involves segmenting the point cloud into distinct spaces and determining the optimal shape for each. This approach, however, struggles with flexible parametric wall geometry and unenclosed spaces. On the other hand, the wall-based approach clusters point cloud data around each wall, aiming to reconstruct topologically consistent structures [19].
The Cloud2BIM’s Identify_walls function, described using the flowchart in Figure 10, detects and defines wall geometry within a point cloud. It isolates individual wall segments within each floor’s horizontal section, set by default to 90–100% of the floor height to minimize noise and disruptive objects. The function begins by filtering points that fall within these height limits and proceeds to extract corresponding $x$- and $y$-coordinates, forming a 2D array (points_2d) that represents the horizontal section of the point cloud at the specified height range.
Figure 10: Flowchart diagram illustrating the algorithms for segmentation of walls. The implementation of this workflow can be found in the aux_functions.py file, function identify_walls, lines 815–970 [32].
This 2D point cloud then undergoes histogram analysis to create a 2D histogram, where each bin’s value reflects the point density in that spatial region. Converting this histogram into a binary image based on a predefined threshold isolates regions with high point densities (likely indicative of walls). This binary image (Figure 11) serves as the basis for further wall detection and geometry definition. It undergoes processing to extract contours using the findContours function from the OpenCV library [42], explicitly targeting the outermost boundaries of wall segments.
@@
As for slabs, each detected contour is approximated into simpler line segments using the Douglas-Peucker algorithm. The choice of epsilon significantly affects the quality of the resulting geometry: a very small epsilon value results in contours with many short segments, leading to an increase in computational time during the subsequent pairing of opposing wall surfaces to form volumetric walls (although these short, colinear segments are merged afterward). Conversely, an excessively large epsilon value may erroneously merge segments that represent separate surfaces, such as distinct walls near openings or shorter wall segments. Therefore, to preserve the geometric accuracy, epsilon must be carefully selected and should never exceed the length of the smallest segment intended to be represented. Our analysis indicates a reasonably wide range of epsilon values within which optimal contour simplification can be reliably achieved (see Figure 12).
Figure 12: Illustration of the Douglas-Peucker algorithm applied to contour approximation, demonstrating the influence of the epsilon parameter on contour simplification and wall surface identification; small epsilon values produce detailed and ragged contours with an excessive number of segments, while large epsilon values overly simplify contours, potentially merging separate surfaces near openings and short segments.